OpenStudy (anonymous):

Simplify the following fraction: ((1/x)-x)/((1/x)+1) I'll draw it too

OpenStudy (anonymous):

(-x+1/x)/[1+1/x]

OpenStudy (mertsj):

\[\frac{\frac{1}{x}-x}{\frac{1}{x}+1}\]

OpenStudy (anonymous):

(1 - x^2 ) / (x + 1) = 1-x

OpenStudy (mertsj):

Is that it?

OpenStudy (anonymous):

[[-x^2+1]/2]/[x+1]/2

OpenStudy (anonymous):

1-x^2/1+x

OpenStudy (anonymous):

final ans

OpenStudy (mertsj):

Multiply numerator and denominator by x and get: \[\frac{1-x ^{2}}{1+x}=\frac{(1-x)(1+x)}{1+x}=1-x\]

OpenStudy (anonymous):



OpenStudy (anonymous):

[1-x][1+x]/[1+x]=1-x

OpenStudy (mertsj):